please help me solve this problem
find three consecutive odd integers such that their sum is five more than 4 times the larges integer

Answer by mananth(16946)

let the integers be x,x+1,x+2
x+x+1+x+2 = 5+4(x+2)
3x+3=5+4x+8
4x-3x=-13+3
x=-10
the numbers are -10,-9,-8